40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 540     Entry time: Wed Jun 18 18:20:10 2008
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: Investigation on the NPRO temperature stabilization glitches 
As Rob pointed out in http://dziban.ligo.caltech.edu:40/40m/537 the MOPA NPRO has been showing some glitchiness in the LTEC loop.
Following Rana's suggestion, Steve and I opened the MOPA and directed a heat gun for a minute to the NPRO hoping that we can see something in the LTEC loop.
The first attachment shows the behavior of LTMP and LTECH along with DTMP and DTECH at the time of the heat gun attack.
T=0 is the time when Steve directed the heat gun to the NPRO. There is no response neither in LTMP nor LTECH.
DTMP and DTECH look like responding.
Around the center, there is a dip in LTMP. This might be caused by removing the heat gun. But we are not sure. This kind of small glitches can be found in LTMP everywhere (see the attachment 2).
It looks like the LTMP sensor is not working, or the LTECH loop is actually working but the LTECH reading is broken.
However, the scan of the slow actuator (temperature) shows the LTECH loop is actually working. So it is a bit confusing.
More investigation is necessary.
See the next entry by me.
Attachment 1: LTEC-loop-HeatGun-Response.pdf  7 kB  | Hide | Hide all
LTEC-loop-HeatGun-Response.pdf
Attachment 2: LTMP-glitches.pdf  30 kB  | Hide | Hide all
LTMP-glitches.pdf
ELOG V3.1.3-